1. what is the probability that a randomly selected mango will weigh less than 1.5 kilos?

1
Expert's answer
2021-12-20T16:03:58-0500

Incomplete question.

Let us take an example related to this.

Q. There are 100 mangoes in a shipment. It is found that it a mean weight of 2.5 kilos with a standard deviation of 2 kilos. what is the probability that a randomly selected mango will weigh less than 1.5 kilos?

A.

Given: Mean "(\mu)=2.5," S.D "(\sigma)=2"

"P(X<1.5)\\

=P(Z<\frac{1.5-2.5}{2})\\

=P(Z<-0.5)\\

=0.3085\\"

Hence, the probability that a randomly selected mango will weigh less than 1.5 kilos is 0.3085.